OpenStudy (anonymous):

plz help i have alot more so please don't get impatient Multiply. Express the product in simplest form. 2 over 6 times 24 equals m A. 1 over 72 B. 8 C. 48 D. 72

OpenStudy (nincompoop):

Use the rule for multiplying fractions \( \large \frac{2}{6} \times \frac{24}{1}= m\)

OpenStudy (nincompoop):

But you can also simplify it first. If you noticed that 24 is divisible by 6

OpenStudy (anonymous):

yes i got the answer but how do i simplify it?

OpenStudy (anonymous):

nvermind

OpenStudy (anonymous):

@bullkidd

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

what answer did u get .. u said u got an answer ..

OpenStudy (anonymous):